OpenStudy (anonymous):

Question 3 The table below shows the values of f(n) for different values of n.

OpenStudy (yanasidlinskiy):

Where's the table?

OpenStudy (anonymous):

 Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

OpenStudy (anonymous):

@YanaSidlinskiy ?

OpenStudy (yanasidlinskiy):

hmm......not really sure about this one....

OpenStudy (anonymous):

@ParthKohli @willsgirl16 @Kainui @HARSH123 @Taylor<3sRin @DollyAcquah @iPwnBunnies @kc_kennylau @surjithayer @surfer1234 @SeaTurtle113 ?

OpenStudy (anonymous):

Is it not the third 1?

OpenStudy (anonymous):

u just took it off

OpenStudy (anonymous):

@d3Xter , do u know?

OpenStudy (anonymous):

I suspect it is, but not to sure. Can't do the recursive thing now. I'll go with the third.

OpenStudy (anonymous):

:( oh well. I hope your right

OpenStudy (anonymous):

Thx d3Xter. u were right :)

OpenStudy (anonymous):

Thx for confirming :) U much welcome :)